OpenStudy (anonymous):

Add. 7 3/10 +6 1/3+ 2 7/10=??? I will fan and medal!!

OpenStudy (anonymous):

@Liv1234

OpenStudy (liv1234):

We have to convert them into fractions first.

OpenStudy (iluvsoccer):

We need to find a common denominator, in this case it can be 30. :)

OpenStudy (anonymous):

okay

OpenStudy (liv1234):

73/10 + 19/3 + 27/10

OpenStudy (liv1234):

Those are your converted fractions.

OpenStudy (liv1234):

In this case we do need a common denominator since it isn't multiplication.

OpenStudy (anonymous):

yeah

OpenStudy (liv1234):

30 is a common denominator as @ILUVSOCCER mentioned.

OpenStudy (anonymous):

yeah i wrote that down

OpenStudy (liv1234):

But to get the denominators to 30, we multiply the numerator by the number we multiplied the denominator by to get 30.

OpenStudy (anonymous):

okay ill do that

OpenStudy (liv1234):

10*3=30 3*10=30 10*3=30 73*3=? 10*19=? 3*27=?

OpenStudy (liv1234):

219 190 81

OpenStudy (liv1234):

219/30+190/30+81/30 =?

OpenStudy (anonymous):

its seven and 3 over 10 not 73 over 10 just fyi

OpenStudy (iluvsoccer):

 This is one, now the others.